# A Finer View of the Parameterized Landscape of Labeled Graph Contractions

### v1

@@ 1 Introduction
This second problem is closely related to the first. As shown in [41], $H$ is a labeled contraction of $G$ if and only if there exists a common contraction of both $G$ and $H$ to a labeled graph $H^{\prime}$ of size at least $|V(H)|$. This equivalence allows complexity results to be transferred between the two problems. The same work provides additional applications and results.
@@ 3.1 Algorithmic Result
  • An introduce edge node $t$ has one child $t^{\prime}$ with $X_{t}=X_{t^{\prime}}$, and is labeled with an edge $\{u,v\}\in E(G)$ such that $\{u,v\}\subseteq X_{t}$. We say that the edge $\{u,v\}$ is introduced at $t$;
@@
Henceforth, we fix a nice tree decomposition $\mathcal{T}$ of $G\cup H$. This choice is crucial because it ensures that for any edge $(u,v)\in E(H)$, there exists some bag $X_{t}$ such that $u,v\in X_{t}$, a condition that is not guaranteed in a tree decomposition of $G$ alone. For each node $t\in\mathcal{T}$, we define $G_{t}$ and $H_{t}$ to be the subgraphs of $G$ and $H$, respectively, induced by the union of all bags in the subtree of $\mathcal{T}$ rooted at $t$.
We now introduce the key definitions for the DP-state at a node $t$ of the tree-decomposition to convey the main ideas of the algorithm. The formal proof follows.
3.1.1 Witness Structures and Functional Representation
Assume that $H$ is a contraction of $G$, and let $\mathcal{W}=\{W_{1},\dots,W_{|V(H)|}\}$ denote a witness structure for this contraction. We represent $\mathcal{W}$ functionally using a pair of auxiliary functions: a parent function and an originator function.
Definition 3.3 (Parent function).
Given a witness structure $\mathcal{W}$, a function $\sigma^{\prime}:V(G)\to V(G)$ is a parent function with respect to $\mathcal{W}$ if there exists a collection of rooted trees $\{T_{i}\}_{i=1}^{k}$, where each $T_{i}$ is a spanning tree of $G[W_{i}]$ rooted at the unique vertex of $W_{i}\cap V(H)$, such that:
  1. 1.
    For all $u\in W_{i}\setminus V(H)$, $\sigma^{\prime}(u)$ is the parent of $u$ in $T_{i}$; and
  2. 2.
    For the unique vertex $u\in W_{i}\cap V(H)$, $\sigma^{\prime}(u)=u$.
Definition 3.4 (Originator function).
Given a witness structure $\mathcal{W}$, a function $\tau^{\prime}:V(G)\to V(H)$ is an originator function with respect to $\mathcal{W}$ if for every $u\in W_{i}$, $\tau^{\prime}(u)$ is the unique vertex in $W_{i}\cap V(H)$.
Note that any valid originator function $\tau^{\prime}$ defines a partition of $V(G)$ into the witness sets $\mathcal{W}$. Also, for any vertex $u\in V(G)$, there exists a bag that contains both $u$ and $\sigma^{\prime}(u)$, but not necessarily $\tau^{\prime}(u)$.
A pair $(\sigma^{\prime},\tau^{\prime})$ is said to be consistent if for every $u\in V(G)$ $\tau^{\prime}(u)=\tau^{\prime}(\sigma^{\prime}(u))$, i.e. $u$ and its parent $\sigma^{\prime}(u)$ belong to the same witness set. In the remainder of the discussion, we assume all parent-originator pairs to be consistent unless stated otherwise. Although $\tau^{\prime}$ can be derived from $\sigma^{\prime}$ by following parent pointers up to a vertex in $V(H)$, we maintain both functions explicitly for algorithmic clarity.
Definition 3.5 (Witness impression).
Every witness structure $\mathcal{W}$ corresponds to pair $(\sigma^{\prime},\tau^{\prime})$, called its witness-impression, where $\sigma^{\prime}$ is a parent function and $\tau^{\prime}$ is the corresponding originator function.
Definition 3.6 (Seed impression).
Let $(\sigma^{\prime},\tau^{\prime})$ be an impression of a witness structure. A pair of partial functions $(\sigma,\tau)$, with $\sigma:V(G_{t})\to V(G_{t})\cup\{\star\}$ and $\tau:V(G_{t})\to V(H_{t})\cup\{\star\}$, is a seed impression of $(\sigma^{\prime},\tau^{\prime})$ at $t$ if:
  • For every vertex $u\in V(G_{t})\setminus X_{t}$, $\sigma(u)=\sigma^{\prime}(u)$ and $\tau(u)=\tau^{\prime}(u)$.
  • For a vertex $u\in X_{t}$, if $\sigma^{\prime}(u)\in V(G_{t})$ then $\sigma(u)=\sigma^{\prime}(u)$ otherwise $\sigma(u)=\star$. The function $\tau$ is derived from $\sigma$ by following parent pointers up to either a vertex in $V(H)$ or $\star$.
Note that $(\sigma,\tau)$ is not merely the restriction of $(\sigma^{\prime},\tau^{\prime})$ to $V(G_{t})$, since a vertex $u$ such that $\sigma^{\prime}(u)\notin V(G_{t})$ might satisfy $\tau^{\prime}(u)\in V(G_{t})$, but we still assign $\tau(u)=\star$. Note that in Figure 1, $\tau^{\prime}(u_{1})\in G_{t}$ but since $\sigma^{\prime}(u_{1})\notin G_{t}$ we assign $\tau(u_{1})=\star$.
We now classify each vertex in the bag $X_{t}$ into one of three types, according to the seed impression $(\sigma,\tau)$:
Definition 3.7 (Types of vertices).
Let $t\in V(\mathcal{T})$ be a node with bag $X_{t}$, and let $(\sigma,\tau)$ be a seed impression at $t$. Then each vertex $u\in X_{t}$ is assigned one of the following types:
  • Requester: $\sigma(u)\in V(G_{t})$ and $\tau(u)\in V(H_{t})$ (i.e., both are defined); $u$ is ready to request adjacency requirements in $H$.
  • Dependent: $\sigma(u)\in V(G_{t})$ but $\tau(u)=\star$; $u$ has been contracted into its parent and depends on the parent node’s decision regarding the partition.
  • Explorer: $\sigma(u)=\star$ (implying $\tau(u)=\star$); $u$ is still exploring who its parent will be.
We classify the pair of requester vertices into the following three parts.
Definition 3.8 (Requester pair classification).
Let $u,v\in X_{t}$ be requester vertices with $\tau(u)=h_{u}$, $\tau(v)=h_{v}$. We say the pair $(u,v)$ is
  • invalid (inv) if $(h_{u},h_{v})\notin E(H)$ but $\tau^{-1}(h_{u}),\tau^{-1}(h_{v})$ are adjacent in $G_{t}$,
  • unsatisfied (unsat) if $(h_{u},h_{v})\in E(H)$ and $\tau^{-1}(h_{u}),\tau^{-1}(h_{v})$ are not adjacent in $G_{t}$, and
  • satisfied (sat), otherwise.
For a vertex $u\in V(G)$, we define the set $\textsf{des}(u)\subseteq V(G)$ of descendants of $u$ with respect to a function $\sigma^{\prime}:V(G)\to V(G)$ as follows:
$\textsf{des}(u)=\left\{v\in V(G)\;\middle|\;\text{there exists an integer }\ell\geq 0\text{ such that }\sigma^{\prime(\ell)}(v)=u\right\},$
where $\sigma^{\prime(0)}(v)=v$, and for $\ell\geq 1$, $\sigma^{\prime(\ell)}(v)=\sigma^{\prime}(\sigma^{\prime(\ell-1)}(v))$. That is, $\textsf{des}(u)$ consists of all vertices $v$ for which repeated applications of $\sigma^{\prime}$ eventually reach $u$, including $u$ itself.
Definition 3.9 (Pseudo-adjacency).
Let $u,v$ be two vertices in $X_{t}$. Then, we say $u$ and $v$ are pseudo-adjacent if sets $\textsf{des}(u)$ and $\textsf{des}(v)$ are adjacent in $G_{t}$.
3.1.2 Signatures and Dynamic Programming States
Suppose the graph $G$ admits an $H$-witness structure $\mathcal{W}$, and let $(\sigma^{\prime},\tau^{\prime})$ be the corresponding witness-impression as specified in Definition 3.5. Let $T$ be a tree decomposition of the graph $G\cup H$, and let $t$ be a node of $T$.
Definition 3.10 (Partial Solution).
We define a partial solution at node $t$ as a tuple $P=(\sigma,\tau,\mathcal{R},\mathcal{A}),$ with the following components:
  • $(\sigma,\tau)$ is a seed impression of the witness-impression $(\sigma^{\prime},\tau^{\prime})$, restricted to the bag $X_{t}$ at $t$, as specified in Definition 3.6,
  • $\mathcal{R}\subseteq R\times R\times\{\textsf{inv},\textsf{sat},\textsf{unsat}\}$ is a set of labeled pairs, where $R\subseteq X_{t}$ denotes the set of requester vertices contained in $X_{t}$ as per Definition 3.7. Each element $(u,v,\ell)\in\mathcal{R}$ represents a label $\ell\in\{\textsf{inv},\textsf{sat},\textsf{unsat}\}$ assigned to the pair $(u,v)$, as per Definition 3.8,
  • $\mathcal{A}\subseteq X_{t}\times X_{t}$ records unordered pairs of vertices in $X_{t}$ that are pseudo-adjacent to each other, as per Definition 3.9.
We now refer to Figure 1 to provide intuition for the above definitions. The figure shows the graph $H$ together with a part of the tree decomposition of $G\cup H$ (we do not explicitly draw $G$). The rounded outlines indicate the witness structure for the contraction, whenever it exists: each group contains exactly one node of $H$, representing the collection of vertices that contract to it. Note that $\sigma(u_{2})=u_{1}$, $\tau(u_{2})=\star$, and $\tau(u_{1})=\sigma(u_{1})=\star$. Since the edge $(u_{3},u_{4})$ is an edge we processed earlier in the DP, we may store $(u_{7},u_{5},\textsf{sat})\in\mathcal{R}$ for $X_{t}$. On the other hand, because the edge $(u_{6},u_{5})$ has not yet been introduced, the edge $(h_{2},h_{3})$ does not appear in the current structure, so we have $(u_{2},u_{5},\textsf{unsat})$. Furthermore, observe that $(u_{8},u_{5})$ and $(u_{7},u_{5})$ belong to $\mathcal{A}$, whereas $(u_{9},u_{5})\notin\mathcal{A}$. In this way, all relevant sets are defined for the dynamic programming.
$u_{8}$$u_{7}$$u_{1}$$h_{2}$$u_{2}$$u_{9}$$u_{5}$$u_{3}$$h_{1}$$u_{4}$$h_{3}$$u_{6}$ $X_{t}$$G_{t}$RestofGRest~of~G
@@ 3.1.2 Signatures and Dynamic Programming States
We now provide a formal definition for the same.
Note that, as it is standard in the case of dynamic programming on tree decomposition, we will not compute all the partial solution at node $t$ but rather just store unique signatures of all the partial solutions at $t$. The signature $S$ of a partial solution $P=(\sigma,\tau,\mathcal{R},\mathcal{A})$ is $(\sigma|_{X_{t}},\tau|_{X_{t}},\mathcal{R},\mathcal{A})$.
States of the Dynamic Programming Algorithms
Definition 3.11 (Signature).
A state of dynamic programming table will be a tuple $S=(\sigma,\tau,\mathcal{R},\mathcal{A})$ such that
  • $\sigma:X_{t}\to X_{t}\cup\{\star\}$ (which will emulate a parent function),
  • $\tau:X_{t}\to V(H_{t})\cup\{\star\}$ (which will emulate an originator function),
  • $\mathcal{R}\subseteq R\times R\times\{\textsf{inv},\textsf{sat},\textsf{unsat}\}$ is a set of labeled pairs representing a label $\ell\in\{\textsf{inv},\textsf{sat},\textsf{unsat}\}$ assigned to the pair $(u,v)$, as per Definition 3.8, and
  • $\mathcal{A}$ is the collection of pair of vertices in $X_{t}$ that are pseudo-adjacent with each other.
The condition on functions ensures that at present $\sigma$ and $\tau$ are valid candidate to be a parent function and an originator function.
For a node $t$ in tree decomposition $T$, we construct a dynamic-programming algorithm that sets the value $\textbf{d}[S,t]$ to True if and only if $S$ is a signature of some partial solution $P$ at node $t$. Otherwise, we set it to False. Initially, all entries are set to False. For any signature $S=(\sigma,\tau,\mathcal{R},\mathcal{A})$, if $\mathcal{R}$ contains an entry of the type $(u,v,\textsf{inv})$ then $\textbf{d}[S,t]$ is set to False and it will never be changed.
3.1.3 Updating Dynamic Programming Tables
We now proceed to describe a recursive algorithm used to update entries in the dynamic programming table for each node type in a nice tree-decomposition.
@@ 3.1.3 Updating Dynamic Programming Tables
Let $t$ be a leaf node. Then $X_{t}=\emptyset$ and the valid partial solution $P$ is empty functions $\sigma,\tau,\mathcal{R}$ since their domain is a subset of $X_{t}$ which is trivially empty. Hence, the only valid signature consists of four empty sets.
@@
If $t$ is a leaf node then set $\textbf{d}[S,t]=\texttt{True}$ for $S=(\emptyset,\emptyset,\emptyset,\emptyset)$, and $\textbf{d}[S^{\prime},t]=\texttt{False}$ for every other tuple $S^{\prime}$.
@@
Let $t$ be an introduce vertex node in the tree decomposition $T$, with child $t^{\prime}$, such that $X_{t}=X_{t^{\prime}}\cup\{x\}$ for some vertex $x\in V(G)\setminus V(G_{t^{\prime}})$. Let $S=(\sigma,\tau,\mathcal{R},\mathcal{A})$ be a signature at node $t$, and let $S^{\prime}=(\sigma^{\prime},\tau^{\prime},\mathcal{R}^{\prime},\mathcal{A}^{\prime})$ be a signature at node $t^{\prime}$.
Definition 3.12.
We say that $S$ and $S^{\prime}$ are introduce-vertex compatible if the following conditions hold:
  • ($C1$)
    For every $y\in X_{t}\setminus\{x\}$, we have $\sigma(y)=\sigma^{\prime}(y)$ and $\tau(y)=\tau^{\prime}(y)$,
  • ($C2$)
    $\mathcal{R}^{\prime}\subseteq\mathcal{R}$ and $\mathcal{A}^{\prime}=\mathcal{A}$.
Let $I(S)$ denote the set of all signatures at $t^{\prime}$ that are introduce-vertex compatible with $S$.
@@
Let $t$ be an introduce vertex node with child $t^{\prime}$, and let $X_{t}=X_{t^{\prime}}\cup\{x\}$ and $x\notin V(H)$. Let $S=(\sigma,\tau,\mathcal{R},\mathcal{A})$ be a signature at node $t$. Assign $\mathbf{d}[S,t]=\texttt{True}$ if and only if the following conditions are satisfied:
@@
    $\sigma(x)=\star$ and $\tau(x)=\star$,
@@
    there exists a signature $S^{\prime}=(\sigma^{\prime},\tau^{\prime},\mathcal{R}^{\prime},\mathcal{A}^{\prime})\in I(S)$ such that $\mathbf{d}[S^{\prime},t^{\prime}]=\texttt{True}$,
  3. (iii)
    $\mathcal{R}=\mathcal{R}^{\prime}$ and $\mathcal{A}=\mathcal{A}^{\prime}$, i.e., no new elements are introduced in either of the sets.

### v2

@@ 1 Introduction
This second problem is closely related to the first. As shown in [41], $H$ is a labeled contraction of $G$ if and only if there exists a common contraction of both $G$ and $H$ to a labeled graph $H^{\prime}$ of size at least $|V(H)|$. This equivalence allows complexity results to be transferred between the two problems. We refer readers to the same work [41] for additional results and applications.
@@ 2 Preliminaries
The same is illustrated on a toy example in Figure 1.
GG12534MM1254
Figure 1: Contraction along edge $(2,3)$
@@
We illustrate the same in Figure 2.
GG1253467HH1237
Figure 2: Shaded regions in $G$ form the witness sets, each containing exactly one representative vertex (circled). Contracting each witness set yields $H$.
@@ 3.1 Algorithmic Result
  • An introduce edge node $t$ has one child $t^{\prime}$ with $X_{t}=X_{t^{\prime}}$, and is labeled with an edge $(u,v)\in E(G)$ such that $\{u,v\}\subseteq X_{t}$. We say edge $(u,v)$ is introduced at $t$.
@@
Henceforth, we fix a nice tree decomposition $\mathcal{T}$ of $G\cup H$. This choice is crucial because it ensures that for any edge $(u,v)\in E(H)$, there exists some bag $X_{t}$ such that $u,v\in X_{t}$, a condition that is not guaranteed in a tree decomposition of $G$ alone. For each node $t\in\mathcal{T}$, we define $G_{t}$ and $H_{t}$ to be the subgraphs of $G$ and $H$, respectively, induced by the union of all bags in the sub-tree of $\mathcal{T}$ rooted at $t$.
$u_{8}$$u_{7}$$u_{1}$$h_{2}$$u_{2}$$u_{9}$$u_{5}$$u_{3}$$h_{1}$$u_{4}$$h_{3}$$u_{6}$ $W_{1}$ $W_{2}$ $W_{3}$$X_{t}$$G_{t}$RestofGRest~of~G
@@
Assume that $H$ is a contraction of $G$, and let $\mathcal{W}=\{W_{1},\dots,W_{|V(H)|}\}$ denote a hypothetical but fixed witness structure for this contraction. Recall that we refer to the unique vertex $h\in W_{i}\cap V(H)$ as the representative vertex of the witness set $W_{i}$.
Definition 3.3 (Active Vertices).
Let $t\in V(\mathcal{T})$ be a node with bag $X_{t}$. A representative vertex $h\in V(H)$ is active at $t$ if its corresponding witness set $W_{h}\in\mathcal{W}$ intersects with $X_{t}$, i.e., $W_{h}\cap X_{t}\neq\emptyset$. The set $A_{t}\subseteq V(H)$ is the collection of all active vertices at $t$.
Observation 3.1.
The number of active vertices $|A_{t}|$ is upper bounded by the size of the bag, $|A_{t}|\leq|X_{t}|$, as each active vertex corresponds to a unique witness set that intersects $X_{t}$. Furthermore, due to the connectivity properties of the witness sets, a representative vertex that becomes inactive in a bag $X_{t}$ (i.e., its witness set $W_{h}$ no longer intersects $X_{t}$) remains inactive in all subsequent bags in the traversal towards the root.
Definition 3.4 (Witness Impression).
For a given witness structure $\mathcal{W}$, the witness impression on the bag $X_{t}$ of node $t$ is the pair $(\mathcal{P},\tau)$, where:
  • $\mathcal{P}$ is the partition of $X_{t}$ induced by the witness sets $\mathcal{W}$:
    $\mathcal{P}=\{W\cap X_{t}\mid W\in\mathcal{W}\text{ and }W\cap X_{t}\neq\emptyset\}.$
  • $\tau:\mathcal{P}\to(X_{t}\cap V(H))\cup\{\uparrow,\downarrow\}$ is a function that maps each part $P\in\mathcal{P}$ to the status of its unique representative vertex $h\in V(H)$:
    $\tau(P)=\begin{cases}h&\text{if }h\in X_{t}\quad\text{(Representative in Bag)}\\ \downarrow&\text{if }h\in V(H_{t})\setminus X_{t}\quad\text{(Representative Forgotten)}\\ \uparrow&\text{if }h\in V(H)\setminus V(H_{t})\quad\text{(Representative Above)}\end{cases}$
Definition 3.5 (Adjacency Sets).
Consider a witness impression $(\mathcal{P},\tau)$ at node $t$. For any pair of distinct partitions $(P_{i},P_{j})\in\binom{\mathcal{P}}{2}$, let $h_{i}$ and $h_{j}$ be their respective representative vertices. We define the following two sets:
  • $\mathcal{A}_{t}$ is the set of pseudo-adjacent pairs $(P_{i},P_{j})$ such that there is an edge in $G$ between $W_{i}\cap V(G_{t})$ and $W_{j}\cap V(G_{t})$, where $W_{i}\cap X_{t}=P_{i}$ and $W_{j}\cap X_{t}=P_{j}$.
  • $\mathcal{R}_{t}$ is the set of anti-adjacent pairs $(P_{i},P_{j})$ such that the representatives $h_{i}$ and $h_{j}$ are non-adjacent in $H$, i.e., $(h_{i},h_{j})\notin E(H)$.
In other words, $\mathcal{A}_{t}$ stores the pair of partitions such that the adjacency between the corresponding witness sets $W_{i}$ and $W_{j}$ is witnessed by some edge in $E(G_{t})$. It might be possible that both endpoints are completely in $V(G_{t})\setminus X_{t}$. Consider an example of witness sets $W_{1}$ and $W_{3}$ in Figure 3. Similarly, $\mathcal{R}_{t}$ stores the pair of partitions that should remain non-adjacent. This information is obvious if active vertices of the corresponding witness sets are in the bag. However, if one of both of them have been forgotten, the pairs in $\mathcal{R}_{t}$ provides the critical information. For example, in Figure 3, $\mathcal{R}_{t}$ contains a pair $(W_{1},W_{2})$ as their active vertices are non-adjacent in $H$.
We are now in a state to describe our dynamic programming algorithm. The dynamic programming algorithm maintains a table $\texttt{dp}[t;S]$, where the signature $S$ at node $t$ is the tuple $(\mathcal{P},\tau,\mathcal{A}_{t},\mathcal{R}_{t})$. We set $\texttt{dp}[t;S]=\texttt{True}$ if and only if there exists an $H$-witness structure $\mathcal{W}$ of $G$ such that $S$ is the witness impression and adjacency sets derived from $\mathcal{W}$ at node $t$.
Definition 3.6 (Valid Signature).
Consider a partition $\mathcal{P}$ of $X_{t}$, a function $\tau:\mathcal{P}\mapsto(X_{t}\cap V(H))\cup\{\uparrow,\downarrow\}$, and two mutually disjoint sets $\mathcal{A}_{t},\mathcal{R}_{t}\subseteq\binom{\mathcal{P}}{2}$. The tuple $S=(\mathcal{P},\tau,\mathcal{A}_{t},\mathcal{R}_{t})$ is an valid signature if all of the following conditions hold:
  1. 1.
    (Cardinality) There is no partition $P_{i}\in\mathcal{P}$ that contains more than one vertex from $V(H)$.
  2. 2.
    (Unique Mapping) There do not exist two distinct partitions $P_{i},P_{j}\in\mathcal{P}$ such that $\tau(P_{i})=\tau(P_{j})\not\in\{\uparrow,\downarrow\}$.
  3. 3.
    (Adjacency Condition) There do not exist two partitions $P_{i},P_{j}$ such that $(P_{i},P_{j})\in\mathcal{R}_{t}$ and $P_{i}$ and $P_{j}$ are adjacent in $G$ (i.e., there is an edge between $P_{i}$ and $P_{j}$).
  4. 4.
    (Internal DP Constraint) There do not exist two partitions $P_{i},P_{j}$ such that $\tau(P_{i})=\uparrow$, $\tau(P_{j})=\downarrow$, and $P_{i}$ and $P_{j}$ are adjacent in $G$.
An invalid signature is defined as a signature that does not satisfy the conditions of a valid signature.
Lemma 3.2.
For an invalid signature $S$, $\texttt{dp}[t;S]=\texttt{False}$.
Proof.
We demonstrate that if a signature $S=(\mathcal{P},\tau,\mathcal{A}_{t},\mathcal{R}_{t})$ satisfies any of the invalidity conditions, it cannot be derived from any valid $H$-witness structure $\mathcal{W}$ of $G$. Thus, the dp entry must be False.
Let $\mathcal{W}$ be an arbitrary $H$-witness structure.
  1. 1.
    (Cardinality Error) If $P_{i}\in\mathcal{P}$ contains distinct vertices $h_{a},h_{b}\in V(H)$, then the witness set $W_{i}$ (of which $P_{i}$ is a fragment) contains multiple representative vertices. This violates the fundamental property of an $H$-witness structure, which requires each $W_{i}\in\mathcal{W}$ to contract to a single representative in $H$.
  2. 2.
    (Unique Mapping Error) If distinct partitions $P_{i},P_{j}\in\mathcal{P}$ map to the same representative $\tau(P_{i})=\tau(P_{j})=h\in X_{t}\cap V(H)$, then both $P_{i}$ and $P_{j}$ must belong to the same witness set $W_{h}$. Since $\mathcal{P}$ is the partition of $X_{t}$ induced by the distinct sets in $\mathcal{W}$, $P_{i}$ and $P_{j}$ cannot be distinct partitions, a contradiction.
  3. 3.
    (Adjacency Contradiction) If $(P_{i},P_{j})\in\mathcal{R}_{t}$, then their representatives $h_{i}$ and $h_{j}$ are non-adjacent in $H$. If, however, $P_{i}$ and $P_{j}$ are adjacent in $G$, the graph contraction rule requires $h_{i}$ and $h_{j}$ to be adjacent in $H$ to cover this edge. This direct contradiction implies no valid contraction exists for this signature.
  4. 4.
    (Internal DP Constraint) If partitions $P_{i}$ ($\tau(P_{i})=\uparrow$) and $P_{j}$ ($\tau(P_{j})=\downarrow$) are adjacent in $G$, the adjacency is between a witness set whose representative $h_{i}$ is yet to be introduced into the subgraph ($h_{i}\in V(H)\setminus V(H_{t})$) and a set whose representative $h_{j}$ has already been forgotten ($h_{j}\in V(H_{t})\setminus X_{t}$). This configuration represents the case when two witness sets are adjacent in $G$ but their representative vertices are not adjacent in $H$. Note that this situation can not be changed with additional vertices or edges. This structural inconsistency rules out a valid DP state.
Therefore, no valid $H$-witness structure can produce an invalid signature $S$. ∎
In rest of the section, whenever we mention a signature, we assume it to be valid according to Definition 3.6. We now present how the dynamic programming table is updated at each type of nodes.
@@
The following lemma is simple to prove.
@@
Let $t$ be a leaf node. The DP entry is $\texttt{dp}[t;S]=\texttt{True}$ if and only if the signature is the empty tuple, $S=(\emptyset,\emptyset,\emptyset,\emptyset)$.
Proof.
The bag is $X_{t}=\emptyset$. Since the partition $\mathcal{P}$ is defined on $X_{t}$, $\mathcal{P}$ must be empty. Consequently, $\tau$ (with domain $\mathcal{P}$) and the adjacency sets $\mathcal{A}_{t}$ and $\mathcal{R}_{t}$ (defined on pairs of $\mathcal{P}$) must all be empty sets. As $G_{t}$ contains no vertices, the only signature consistent with any partial witness structure is the empty one. ∎
@@
Let $t$ be an introduce vertex node in the tree decomposition $T$, with child $t^{\prime}$, such that $X_{t}=X_{t^{\prime}}\cup\{x\}$ for some vertex $x\in V(G)\setminus V(G_{t^{\prime}})$. Note that as we treat introduce edge node separately, at this stage, we do not have to consider any new edge. Let $S=(\mathcal{P},\tau,\mathcal{A}_{t},\mathcal{R}_{t})$ be a signature at node $t$, and let $S^{\prime}=(\mathcal{P}^{\prime},\tau^{\prime},\mathcal{A}_{t^{\prime}},\mathcal{R}_{t^{\prime}})$ be a signature at node $t^{\prime}$.
Definition 3.7 (Introduce-Vertex Compatibility).
Valid signatures $S=(\mathcal{P},\tau,\mathcal{A}_{t},\mathcal{R}_{t})$ at node $t$ and $S^{\prime}=(\mathcal{P}^{\prime},\tau^{\prime},\mathcal{A}_{t^{\prime}},\mathcal{R}_{t^{\prime}})$ at its child $t^{\prime}$ are introduce-vertex compatible if there exists a mapping $\phi:\mathcal{P}^{\prime}\to\mathcal{P}$ satisfying one of the following two cases:
  1. (i)
    $x$ joins an existing witness set: There exists a unique $P^{\prime}_{i}\in\mathcal{P}^{\prime}$ such that $\phi(P^{\prime}_{i})=P^{\prime}_{i}\cup\{x\}$, and for all other $P^{\prime}_{j}\in\mathcal{P}^{\prime}$, $\phi(P^{\prime}_{j})=P^{\prime}_{j}$.
    • (C1)
      If $x\in V(H)$, then $\tau(\phi(P^{\prime}_{i}))=x$ and $\tau^{\prime}(P^{\prime}_{i})=\uparrow$. Otherwise, $\tau(\phi(P^{\prime}_{j}))=\tau^{\prime}(P^{\prime}_{j})$ for all $P^{\prime}_{j}\in\mathcal{P}^{\prime}$.
    • (C2)
      $\mathcal{A}_{t}=\{(\phi(P^{\prime}_{a}),\phi(P^{\prime}_{b}))\mid(P^{\prime}_{a},P^{\prime}_{b})\in\mathcal{A}_{t^{\prime}}\}$.
    • (C3)
      $\mathcal{R}_{t}=\{(\phi(P^{\prime}_{a}),\phi(P^{\prime}_{b}))\mid(P^{\prime}_{a},P^{\prime}_{b})\in\mathcal{R}_{t^{\prime}}\}$.
  2. (ii)
    $x$ starts a new witness set: $\mathcal{P}=\mathcal{P}^{\prime}\cup\{P_{x}\}$ where $P_{x}=\{x\}$. The mapping $\phi$ is the identity $\phi(P^{\prime})=P^{\prime}$ for all $P^{\prime}\in\mathcal{P}^{\prime}$.
    • (C1)
      If $x\in V(H)$, then $\tau(P_{x})=x$. Otherwise, $\tau(P_{x})=\uparrow$.
    • (C2)
      $\mathcal{A}_{t}=\mathcal{A}_{t^{\prime}}$. (Note that $P_{x}$ cannot be part of any pseudo-adjacency as no edges are introduced).
    • (C3)
      $\mathcal{R}_{t}=\mathcal{R}_{t^{\prime}}\cup\{(P_{x},P_{j})\mid P_{j}\in\mathcal{P}^{\prime}\text{ and }\tau(P_{x}),\tau(P_{j})\in V(H)\text{ and }(\tau(P_{x}),\tau(P_{j}))\notin E(H)\}$.
@@
Let $t$ be an introduce vertex node with child $t^{\prime}$. For a valid signature $S$, $\texttt{dp}[t;S]=\texttt{True}$ if and only if there exists a signature $S^{\prime}$ such that $\texttt{dp}[t^{\prime};S^{\prime}]=\texttt{True}$ and $S,S^{\prime}$ are introduce-vertex compatible.
Proof.
$(\Rightarrow)$ If $\texttt{dp}[t;S]=\texttt{True}$ with witness structure $\mathcal{W}$, its restriction $\mathcal{W}^{\prime}$ to $G_{t^{\prime}}$ induces a signature $S^{\prime}$. Since $x$ has no incident edges in $G_{t}$, the adjacency set $\mathcal{A}_{t}$ is simply the projection of $\mathcal{A}_{t^{\prime}}$. If $x$ joins an existing part $P^{\prime}_{i}$ (Case i), the mapping $\phi$ updates the part. If $x$ starts a new part (Case ii), $\mathcal{R}_{t}$ is correctly initialized to reflect non-adjacency requirements in $H$. In both cases, $S$ and $S^{\prime}$ are compatible.
$(\Leftarrow)$ Given a witness structure for $G_{t^{\prime}}$ inducing $S^{\prime}$, we extend it by adding $x$ to a witness set as specified by $\phi$. Since $x$ is not incident to any edges in $G_{t}$, no new pseudo-adjacencies are created, making the update $\mathcal{A}_{t}=\mathcal{A}_{t^{\prime}}$ consistent. The updates to $\tau$ and $\mathcal{R}_{t}$ ensure that representative identities and non-adjacency requirements are correctly tracked. Thus, the extended structure is a valid partial $H$-witness structure for $G_{t}$. ∎
Introduce Edge Node
Let $t$ be an introduce edge node in the tree decomposition $T$, with child $t^{\prime}$, such that $X_{t}=X_{t^{\prime}}$, and an edge $e=(u,v)\in E(G)$ is introduced at $t$, where $u,v\in X_{t}$. Let $P_{u},P_{v}\in\mathcal{P}$ be the parts containing $u$ and $v$, respectively.
Let $S=(\mathcal{P},\tau,\mathcal{A}_{t},\mathcal{R}_{t})$ be a signature at node $t$, and $S^{\prime}=(\mathcal{P}^{\prime},\tau^{\prime},\mathcal{A}_{t^{\prime}},\mathcal{R}_{t^{\prime}})$ be a signature at node $t^{\prime}$.
Definition 3.8 (Introduce-Edge Compatibility).
Valid signatures $S=(\mathcal{P},\tau,\mathcal{A}_{t},\mathcal{R}_{t})$ at node $t$ and $S^{\prime}=(\mathcal{P}^{\prime},\tau^{\prime},\mathcal{A}_{t^{\prime}},\mathcal{R}_{t^{\prime}})$ at child $t^{\prime}$ are introduce-edge compatible if all of the following conditions hold:
  • (C1)
    $\mathcal{P}=\mathcal{P}^{\prime}$, $\tau=\tau^{\prime}$, and $\mathcal{R}_{t}=\mathcal{R}_{t^{\prime}}$.
  • (C2)
    Adjacency Update:
@@
      If $P_{u}=P_{v}$, then $\mathcal{A}_{t}=\mathcal{A}_{t^{\prime}}$.
@@
      If $P_{u}\neq P_{v}$, then $\mathcal{A}_{t}=\mathcal{A}_{t^{\prime}}\cup\{(P_{u},P_{v})\}$.
Lemma 3.5.
Let $t$ be an introduce edge node with child $t^{\prime}$, and let $e=(u,v)$ be the introduced edge. Then, $\texttt{dp}[t;S]=\texttt{True}$ if and only if there exists a signature $S^{\prime}\in\mathcal{S}_{t^{\prime}}$ such that $\texttt{dp}[t^{\prime};S^{\prime}]=\texttt{True}$ and $S$ and $S^{\prime}$ are introduce-edge compatible.